# Understanding User Mental Models in AI-Driven Code Completion Tools: Insights from an Elicitation Study

### v1

@@ 6 Prototype Development
  • Suggestion Length: A slider ranging from 1 to 5 adjusts the granularity of suggestions, balancing concise completions with detailed recommendations tailored to the user’s needs.
@@
Moreover, ATHENA allows the developer to receive explanations for selected pieces of code within the editor via contextual IDE options. Specifically, a light-bulb icon appears at the top of the selected code, allowing for three distinct actions:
@@
Regardless of the solutions used by the developer to invoke the code suggestions and explanations, ATHENA’s output can be visualized in two ways, according to the developer’s preferences. The first solution consists of showing code/explanations within the code editor (either inline or in a tooltip). Another option is to show suggestions and explanations separately from the code editor within a side panel, as shown in Figure 5:
Figure 5: Prototype interface view in case the user prefers suggestions to be delivered in a separate panel.
@@ 6.1 Use Case Scenarios
In the following, two scenarios are reported to clarify how the ATHENA tool can be installed and used the first time (Scenario 1, Section 6.1.1) and how the tool can be used by the same users in different contexts, thus demonstrating its suitability to different mental models of the same users (Scenario 2, Section 6.1.2). Two videos of the two scenarios are also attached to show how ATHENA works.
@@ 6.1.2 Scenario 2: Adapting to Debugging Needs
During debugging, Alex selects a block of code that seems problematic. A lightbulb icon appears next to the selected code, offering three options: “Explain the Code (Concise)”, “Explain the Code (Detailed)”, and “Open Chatbot”. Alex selects the first option (Explain the Code (Concise)) and promptly receives a brief summary of the code’s functionality. However, Alex wants to delve deeper into the details of the generated code; therefore, they select the second option (Explain the Code (Detailed)), which provides a more in-depth explanation, helping them understand the logic and potential pitfalls. Finally, to receive additional guidance, Alex selects the Open Chatbot option; this invokes the chatbot, which provides context-aware suggestions for refactoring the code.
@@
By tailoring the tool’s behavior to their specific debugging needs, Alex maintains focus and leverages the extension’s features to complete the task effectively. The combination of personalized settings, code actions, and a context-aware chatbot reinforces the tool’s versatility and value in diverse workflows.
@@ 7 Conclusion
Limitations of this study mostly regard the convenience sampling methodology that was adopted, which led to a sample of very young participants, with limited to no expertise in companies. Therefore, despite the large sample size, the findings of this study are not directly generalizable to professional developers. To address this limitation we plan to conduct usablity studies with different categories of users. Such studies will be initially conducted in a laboratory setting, and then in-the-wild within an IT company; this will allow us to measure the effectiveness and usability of ATHENA compared to other popular CCTs. Moreover, we will observe how different features of a CCT can affect the usability and performance of users with varying degrees of expertise with different categories of users. Future work will also entail further developing our CCT and extending its functionalities by implementing the remaining features that emerged from the elicitation study. Moreover, we will explore the use of different LLMs as the AI code completion engine of the tool; in fact, to generate code suggestions and explanations ATHENA employs OpenAI’s GPT-4o mini, an LLM which has a limited size but performs well on coding problems [48]. However, different models may prove more suitable for use in corporate settings. For instance, a non-commercial solution such as Meta’s open-source LLM LLama [49] could be deployed on a company’s server to prevent sharing confidential data, such as project repositories, with third parties. Furthermore, models of varying sizes could be contemplated in light of a tradeoff between costs and performance.
@@ Acknowledgements
Authors Contribution Statement
@@ Appendix A Themes and Codes
Table 2: Themes for RQ1
RQ1. Where to show the code completion in the UI?
@@
Separate window 6 Code 1: ”Generated code should appear in a side window or tab” sideWindow: 1,3,4,5 4
Code 2: ”Generated code should appear in a tab at the bottom of the IDE” bottomWindow: 3 2
Code 3: ”Generated code should appear in a side movable window” movableTab: 7 2
Popup 3 Code 1: ”Pop-up near the main code” 2
Code 2: ”Toggable pop-ups to view code completions” 1
@@
Adaptive visualization 2 Code 1: ”Separate area for complex hints, inline for short ones” 2
Table 3: Themes for RQ2
RQ2. When should the suggestion be activated during the interaction?
@@
Proactive 4 Code 1: ”Trigger hint during typing automatically” 2
Code 2: ”Hints appear when user is blocked” 4
Manual 5 Code 1: ”Manual trigger with shortcut” 2
Code 2: ”Manual activation preferred over proactive activation” 2
Code 3: ”Manual activation by selecting code” 2
Code 4: ”Manual activation option required” 3
Code 5: ”On-demand activation via comment” 1
Hybrid 6 Code 1: ”Automatic hints during typing, manual by selection” 5
Code 2: ”Hybrid: automatic for obvious hints, manual otherwise” 1
Code 3: ”Manual hints for user when blocked” 1
When coding context is available 4 Code 1: ”Hints appear after analyzing typing context” 4
Table 4: Themes for RQ3
RQ3. How should the suggested code be shown?
@@
Inline 5 Code 1: ”Change font to transparent for generated code” 3
@@
Code 4: ”Color-coded suggestion types” 1
Code 5: ”Toggle color-coding for AI/human code” 1
Separate area 7 Code 1: ”Separate window for detailed suggestions” 6
Code 2: ”Separate place for code suggestions” 1
@@
Clicking an icon or badge 2 Code 1: ”Pressable icon to signal suggestions availability” 2
Chatbot 3 Code 1: ”The code suggestion should happen as a dialogue with a chatbot” 2
@@
Table 5: Themes for RQ4
RQ4. What should be suggested?
@@
Controlling granularity 7 Code 1: ”Code should be given at varying granularity, choosable by the user” 2
Code 2: ”Code should be generated at a scope level defined by the user” 4
Code 3: ”Generated code should be limited to the function scope level” 1
Code 4: ”Generated code should be limited to the file scope level” 2
Code 5: ”Generated code should have a minimum amount of tokens” 1
Code 6: ”Generated code should at least be one line of code” 1
Different alternatives offered 3 Code 1: ”Multiple code alternatives to choose from are preferred” 2
Code 2: ”In certain contexts, users should be able to choose from multiple code alternatives” 2
Minimal suggestions, details on-demand 4 Code 1: ”Code should not be given right away, but be shown as an expandable preview” 1
Code 2: ”Suggested code should be given incrementally to foster the user’s reasoning” 3
Code 3: ”Single suggestions (with no alternatives) are preferred” 1
Code 4: ”The code completion could be limited to just one token” 1
Codes and documentation 2 Code 1: ”Show documentation together with the suggested code” 2
Code adapted to the context 3 Code 1: ”The generated code should adapt to the context of the user (user profile and their work environment)” 3
Table 6: Themes for RQ5
RQ5. How should the suggestion be customized?
@@
Activation timing of suggestions and explanations 6 Code 1: ”Adjust activation timing manually” 2
Code 2: ”Enable/Disable automatic completion” 3
Code 3: ”Customize activation shortcut” 1
Code 4: ”Customize position of toggle button” 1
@@
Appearance 6 Code 1: ”Adjust font and style of suggestions” 4
Code 2: ”Define the position of the suggestions” 4
Coding style 4 Code 1: ”Set the coding style for the generated code” 3
Code 2: ”Set the standards to comply with for the generated code” 3
Code 3: ”Define which data is used for fine tuning the code generation model” 2
Granularity of suggestions and explanations 6 Code 1: Customize the explanation granularity 2
Code 2: ”Control the presence of comments in the generated code” 2
Code 3: ”Control amount of code provided by tool” 4
Code 4: ”Number of different suggestions given” 1
Code 5: ”Limit the scope of the code generation” 2
Table 7: Themes for RQ6
RQ6. Where to show the code completion explanation in the UI?
@@
Inline comments 6 Code 1: ”Comment embedded in the code for context” 4
Code 2: ”Explanation be shown together with the generated code” 2
Code 3: ”Avoid to show complete explanations in comments but only show a brief title” 1
Code 4: ”Display explanation next to the code suggestion” 1
Separate window 7 Code 1: ”Separate window or tab for explanations” 3
Code 2: ”Define the position of the suggestions” 1
@@
Code 4: ”Display explanation in a dedicated lateral window” 4
Code 5: ”Show detailed explanation in a separated window” 1
Code 6: ”Provide explanation in a distinct popup” 2
Code 7: ”Sidebar tab to show explanations alongside code completion” 3
Chatbot 3 Code 1: ”Explanations can be asked to a chatbot” 3
Table 8: Themes for RQ7
RQ7. When to show the explanation during the interaction?
@@
Manually interacting with the suggested code 6 Code 1: ”Explanations should be provided for pieces of code selected by the user” 3
Code 2: ”Explanations appear when clicking the mouse” 2
Code 3: ”Explanations appear when hovering over generated code” 2
Code 4: ”Explanations should be manually expandable” 3
Shortcut 3 Code 2: ”Explanations are invoked by clicking a key or combination of keys” 3
Automatic short explanation, details on-demand 3 Code 1: ”Automatic explanation for short suggestions” 3
Code 2: ”Explanation shown automatically to non-expert users” 1
Code 3: ”Long explanations should be provided only on demand” 1
Table 9: Themes for RQ8
RQ8. What should be explained?
@@
Purpose 5 Code 1: ”Explain why the code was generated in a specific manner” 4
Code 2: ”List pros and cons about the generated code” 1
Code 3: ”Explain different possible alternatives” 1
Examples and references 5 Code 1: ”Provide real-world examples for better understanding” 4
Code 2: ”Add documentation references” 3
Code 3: ”Use trusted sources to report additional information in the explanation” 2
Contextual explanation 4 Code 1: ”Provide suggestion of the selected piece of code” 4
Minimal explanation, details on-demand 4 Code 1: ”Explanation should be minimal, with the possibility to manually ask for more detail” 2
Code 2: ”It would be useful to provide the explanation as a dialogue in a chat interaction” 1
Code 3: ”Provide an explanation preview that is expandable onDemand” 1
Functionality 4 Code 1: ”Explanation should describe what the code is and does” 3
Code 2: ”Explanation should cover what data structures and code constructs appear in the generated code” 2
Code 3: ”Show pseudocode or flowcharts” 2
Code 4: ”Explanation should cover aspects that are not familiar to the user” 1
Code 5: ”Explanation should also explain low level concepts such as syntax and data types” 1

### v2

@@ 7 Reifying the mental model: the ATHENA prototype
  • Suggestion Length: A drop-down menu with options ranging from 1 to 10 adjusts the granularity of suggestions, balancing concise completions with detailed recommendations tailored to the user’s needs.
@@
ATHENA allows the developer to receive explanations for selected pieces of code within the editor via contextual IDE options. Specifically, a light-bulb icon appears at the top of the selected code, allowing for three distinct actions:
@@
Regardless of the solutions used by the developer to invoke the code suggestions and explanations, ATHENA’s output can be visualized in two ways, according to the developer’s preferences. The first solution consists of showing code/explanations within the code editor (either inline or in a tooltip). Another option is to show suggestions and explanations separately from the code editor within a side panel, as shown in Figure 6.
Figure 6: ATHENA interface: the user prefers to show suggestions in a separate panel.
@@ 7.1 Use case scenarios
In the following, two scenarios are reported to clarify how the ATHENA tool can be installed and used the first time (Scenario 1, Section˜7.1.1) and how the tool can be used by the same users in different contexts, thus demonstrating its suitability to different mental models of the same users (Scenario 2, Section˜7.1.2). Two videos of the two scenarios are also attached to show how ATHENA works.
@@ 7.1.2 Scenario 2: adapting to debugging needs
During debugging, Alex selects a block of code that seems problematic. A lightbulb icon appears next to the selected code, offering three options: “Explain the Code (Concise)”, “Explain the Code (Detailed)”, and “Open Chatbot”. Alex selects the first option (Explain the Code (Concise)) and promptly receives a brief summary of the code’s functionality. Alex wants to delve deeper into the details of the generated code; therefore, Alex selects the second option (Explain the Code (Detailed)), which provides a more in-depth explanation, to help the people understand the logic and potential pitfalls. To receive further guidance, Alex selects the Open Chatbot option; this invokes the chatbot, which provides context-aware suggestions for refactoring the code.
@@
By tailoring the tool’s behavior to their specific debugging needs, Alex maintains the focus and leverages the extension’s features to complete the task effectively. The combination of personalized settings, code actions, and a context-aware chatbot reinforces the tool’s versatility and value in diverse workflows.
8 Threats to validity
Even if our study provides useful insights about developers’ mental models on CCTs, several limitations must be acknowledged.
External validity
The participants involved in our study were recruited among the students of two Italian universities, primarily third-year undergraduate computer science students. Although this sampling is appropriate for our exploratory goals, it does not represent the broader population of professional developers. Moreover, the sample skewed heavily male (48 out of 56), and few participants had extensive industry experience. In addition, students may lack the experience necessary to generate highly innovative or realistic design expectations. As a consequence, the generalizability of the findings to different developer populations is limited. Future research is needed to validate and extend these insights through studies involving participants from different industry settings and different skills, with more females, coming from multiple countries, and with more varied demographic backgrounds.
Internal validity
The workshop methodology adopted in our study, while useful for eliciting user mental models, can introduce biases. For instance, some participants may dominate the discussion while others may be quieter and less participatory. To mitigate this bias, we structured the workshops introducing an individual reflection phase for each question, and encouraged all participants to voice their opinions before group synthesis. Another potential problem regards the fact that researcher facilitation may still have unintentionally shaped some discussions. To mitigate this bias, audio recordings have been triangulated with observer transcription before performing the qualitative analysis.
Construct validity
The eight questions used in the workshops to guide the discussion were derived from the analysis of the existing literature on CCTs, with the aim of covering the critical interaction dimensions of CCTs. This structure supported a systematic coverage of relevant aspects; however, it may have constrained participants from reporting issues outside the predefined scope. To mitigate this problem, we introduced open-ended discussion and sketching tools that helped surface unanticipated questions.
Methodological limitations
Co-design workshops tend to foster breadth and richness of group discussion. However, participants, in particular students, may find it difficult to anticipate how they would interact with a real CCT, thus introducing a form of hypothetical bias. This is particularly critical for early-stage technologies, as in the case of our study. Furthermore, participants’ comments may not fully reflect their behavior in the real world. As future work, to mitigate this problem, complementary methods should be adopted, for example, individual interviews, in-the-wild deployment, and usage logging.
@@ 9 Conclusion
Limitations of this study mostly regard the convenience sampling methodology that was adopted, which led to a sample of very young participants, with limited to no expertise in companies. Therefore, despite the large sample size, the findings of this study are not directly generalizable to professional developers. To address this limitation we plan to conduct usablity studies with different categories of users. Such studies will be initially conducted in a laboratory setting, and then in-the-wild within an IT company; this will allow us to measure the effectiveness and usability of ATHENA compared to other popular CCTs. Moreover, we will observe how different features of a CCT can affect the usability and performance of users with varying degrees of expertise with different categories of users. Future work will also entail further developing our CCT and extending its functionalities by implementing the remaining features that emerged from the elicitation study. Moreover, we will explore the use of different LLMs as the AI code completion engine of the tool; in fact, to generate code suggestions and explanations ATHENA employs OpenAI’s GPT-4o mini, an LLM which has a limited size but performs well on coding problems (OpenAI, 2024b). However, different models may prove more suitable for use in corporate settings. For instance, a non-commercial solution such as Meta’s open-source LLM LLama (Meta, 2024) could be deployed on a company’s server to prevent sharing confidential data, such as project repositories, with third parties. Furthermore, models of varying sizes could be contemplated in light of a tradeoff between costs and performance.
@@ Acknowledgements
CRediT authorship contribution statement
@@ Appendix A Themes and codes
Table 2: Themes for Question 1
Q1. Where to show the code completion in the UI?
@@
Separate window 6 Code 1: "Generated code should appear in a side window or tab" sideWindow: 1,3,4,5 4
Code 2: "Generated code should appear in a tab at the bottom of the IDE" bottomWindow: 3 2
Code 3: "Generated code should appear in a side movable window" movableTab: 7 2
Popup 3 Code 1: "Pop-up near the main code" 2
Code 2: "Toggable pop-ups to view code completions" 1
@@
Adaptive visualization 2 Code 1: "Separate area for complex hints, inline for short ones" 2
Table 3: Themes for Question 2
Q2. When should the suggestion be activated during the interaction?
@@
Proactive 4 Code 1: "Trigger hint during typing automatically" 2
Code 2: "Hints appear when user is blocked" 4
Manual 5 Code 1: "Manual trigger with shortcut" 2
Code 2: "Manual activation preferred over proactive activation" 2
Code 3: "Manual activation by selecting code" 2
Code 4: "Manual activation option required" 3
Code 5: "On-demand activation via comment" 1
Hybrid 6 Code 1: "Automatic hints during typing, manual by selection" 5
Code 2: "Hybrid: automatic for obvious hints, manual otherwise" 1
Code 3: "Manual hints for user when blocked" 1
When coding context is available 4 Code 1: "Hints appear after analyzing typing context" 4
Table 4: Themes for Question 3
Q3. How should the suggested code be shown?
@@
Inline 5 Code 1: "Change font to transparent for generated code" 3
@@
Code 4: "Color-coded suggestion types" 1
Code 5: "Toggle color-coding for AI/human code" 1
Separate area 7 Code 1: "Separate window for detailed suggestions" 6
Code 2: "Separate place for code suggestions" 1
@@
Clicking an icon or badge 2 Code 1: "Pressable icon to signal suggestions availability" 2
Chatbot 3 Code 1: "The code suggestion should happen as a dialogue with a chatbot" 2
@@
Table 5: Themes for Question 4
Q4. What should be suggested?
@@
Controlling granularity 7 Code 1: "Code should be given at varying granularity, choosable by the user" 2
Code 2: "Code should be generated at a scope level defined by the user" 4
Code 3: "Generated code should be limited to the function scope level" 1
Code 4: "Generated code should be limited to the file scope level" 2
Code 5: "Generated code should have a minimum amount of tokens" 1
Code 6: "Generated code should at least be one line of code" 1
Different alternatives offered 3 Code 1: "Multiple code alternatives to choose from are preferred" 2
Code 2: "In certain contexts, users should be able to choose from multiple code alternatives" 2
Minimal suggestions,
details on-demand
4 Code 1: "Code should not be given right away, but be shown as an expandable preview" 1
Code 2: "Suggested code should be given incrementally to foster the user’s reasoning" 3
Code 3: "Single suggestions (with no alternatives) are preferred" 1
Code 4: "The code completion could be limited to just one token" 1
Codes and documentation 2 Code 1: "Show documentation together with the suggested code" 2
Code adapted to the context 3 Code 1: "The generated code should adapt to the context of the user (user profile and their work environment)" 3
Table 6: Themes for Question 5
Q5. How should the suggestion be customized?
@@ 4 Code 1: "Code should not be given right away, but be shown as an expandable preview" 1
Activation timing of
suggestions and explanations
6 Code 1: "Adjust activation timing manually" 2
Code 2: "Enable/Disable automatic completion" 3
Code 3: "Customize activation shortcut" 1
Code 4: "Customize position of toggle button" 1
@@ 6 Code 1: "Adjust activation timing manually" 2
Appearance 6 Code 1: "Adjust font and style of suggestions" 4
Code 2: "Define the position of the suggestions" 4
Coding style 4 Code 1: "Set the coding style for the generated code" 3
Code 2: "Set the standards to comply with for the generated code" 3
Code 3: "Define which data is used for fine tuning the code generation model" 2
Granularity of suggestions
and explanations
6 Code 1: Customize the explanation granularity 2
Code 2: "Control the presence of comments in the generated code" 2
Code 3: "Control amount of code provided by tool" 4
Code 4: "Number of different suggestions given" 1
Code 5: "Limit the scope of the code generation" 2
Table 7: Themes for Question 6
Q6. Where to show the code completion explanation in the UI?
@@ 6 Code 1: Customize the explanation granularity 2
Inline comments 6 Code 1: "Comment embedded in the code for context" 4
Code 2: "Explanation be shown together with the generated code" 2
Code 3: "Avoid to show complete explanations in comments but only show a brief title" 1
Code 4: "Display explanation next to the code suggestion" 1
Separate window 7 Code 1: "Separate window or tab for explanations" 3
Code 2: "Define the position of the suggestions" 1
@@
Code 4: "Display explanation in a dedicated lateral window" 4
Code 5: "Show detailed explanation in a separated window" 1
Code 6: "Provide explanation in a distinct popup" 2
Code 7: "Sidebar tab to show explanations alongside code completion" 3
Chatbot 3 Code 1: "Explanations can be asked to a chatbot" 3
Table 8: Themes for Question 7
Q7. When to show the explanation during the interaction?
@@
Manually interacting with
the suggested code
6 Code 1: "Explanations should be provided for pieces of code selected by the user" 3
Code 2: "Explanations appear when clicking the mouse" 2
Code 3: "Explanations appear when hovering over generated code" 2
Code 4: "Explanations should be manually expandable" 3
Shortcut 3 Code 2: "Explanations are invoked by clicking a key or combination of keys" 3
Automatic short explanation,
details on-demand
3 Code 1: "Automatic explanation for short suggestions" 3
Code 2: "Explanation shown automatically to non-expert users" 1
Code 3: "Long explanations should be provided only on demand" 1
Table 9: Themes for Question 8
Q8. What should be explained?
@@ 3 Code 1: "Automatic explanation for short suggestions" 3
Purpose 5 Code 1: "Explain why the code was generated in a specific manner" 4
Code 2: "List pros and cons about the generated code" 1
Code 3: "Explain different possible alternatives" 1
Examples and references 5 Code 1: "Provide real-world examples for better understanding" 4
Code 2: "Add documentation references" 3
Code 3: "Use trusted sources to report additional information in the explanation" 2
Contextual explanation 4 Code 1: "Provide suggestion of the selected piece of code" 4
Minimal explanation,
details on-demand
4 Code 1: "Explanation should be minimal, with the possibility to manually ask for more detail" 2
Code 2: "It would be useful to provide the explanation as a dialogue in a chat interaction" 1
Code 3: "Provide an explanation preview that is expandable onDemand" 1
Functionality 4 Code 1: "Explanation should describe what the code is and does" 3
Code 2: "Explanation should cover what data structures and code constructs appear in the generated code" 2
Code 3: "Show pseudocode or flowcharts" 2
Code 4: "Explanation should cover aspects that are not familiar to the user" 1
Code 5: "Explanation should also explain low level concepts such as syntax and data types" 1